can you solve using substitution 7x+11y=188 13x+11y=236

Answer by greenestamps(13200)



Yes, I can....

But that's probably not what you meant.

The term "11y" is in both of the given equations. So

7x+11y=188 --> 11y=188-7x
13x+11y=236 --> 11y=236-13x

Then, by substitution,

188-7x=236-13x

From there, solve using basic algebra.